OpenStudy (anonymous):

find the slope: x/7+y/5=2

OpenStudy (amistre64):

slope = -a/b in this case: -1/7 ------ 1/5

OpenStudy (amistre64):

which of course simplifies to what ever :)

OpenStudy (anonymous):

5x+7y=70 ===> y=(-5/7)x+10

OpenStudy (anonymous):

slope = -5/7

OpenStudy (anonymous):

thank you!!